The hot water tap of a bath delivers water at 80°c at a rate of 10kg min-1.The cold water tap of the bath delivers water at 20°c at the rate of 20kg min-1. Assuming that both taps are left for 5minutes, calculate the final temperature of the bath water


1
Expert's answer
2021-05-08T14:39:33-0400

Solution.

"t_1=80^oC;"

"v_1=10kg\/min;"

"t_2=20^oC;"

"v_2=20kg\/min;"

"\tau=5min;"

"t-?;"

"m_1=v_1\tau;"

"m_1=10kg/min\sdot5min=50kg;"

"m_2=v_2\tau;"

"m_2=20kg/min\sdot5min=100kg;"

"Q=cm\Delta t;"

"cm_1(t_1-t)=cm_2(t-t_2)\implies t=\dfrac{m_2t_2+m_1t_1}{m_2+m_1};"

"t=\dfrac{100kg\sdot20^oC+50kg\sdot80^oC}{100kg+50kg}=40^oC;"


Answer: "t=40^oC."